OpenStudy (anonymous):

Solve. One point for showing the steps to solve each equation and one point for the correct solution. 2(n - 3) + 5 = 3(n - 1) Solve. One point for showing the steps to solve each equation and one point for the correct solution. 2(n - 3) + 5 = 3(n - 1) @Mathematics

OpenStudy (sasogeek):

before i write the solution, just so u know, it's not necessary to tell us what points u get for answering it. besides, you'r enot going to give me any points here on this site for solving it

OpenStudy (sasogeek):

2(n-3) + 5 = 3(n-1) 2n-6+5=3n-3 -6+5+3=3n-2n 2=n =>n=2